### v3

# Principal stratification with $U$-statistics under principal ignorability

Abstract
Principal stratification is a popular framework for causal inference in the presence of an intermediate outcome. While the principal average treatment effects have traditionally been the default target of inference, it may not be sufficient when the interest lies in the relative favorability of one potential outcome over the other within the principal stratum. We thus introduce the principal generalized causal effect estimands, which extend the principal average causal effects to accommodate nonlinear contrast functions. Under principal ignorability, we expand the theoretical results in Jiang et al. (2022) to a much wider class of causal estimands in the presence of a binary intermediate variable. We develop identification formulas and derive the efficient influence functions of the generalized estimands for principal stratification analyses. These efficient influence functions motivate a set of multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on $U$-statistics. The proposed methods are illustrated through simulations and the analysis of a data example.
Keywords: Causal inference, efficient influence function, principal stratification, multiply robust estimation, $U$-statistics, win ratio
1 Introduction
For causal inference under the Neyman-Rubin model (Rubin, 1974), the average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest. Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$. Conceptually, the average causal effect estimand can be expanded to a wider class, allowing arbitrary comparisons. That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$. Examples include the average causal effect when $h(u,v)=u-v$, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$ (Lu et al., 2020). Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$. However, a closely related quantity, the generalized causal effect (O’brien, 1988; Mao, 2017), can be formulated as:
$\tau_{h}=\int\int h(u,v)\mathrm{d}F_{1}(u)\mathrm{d}F_{0}(v),$ (1)
where $F_{z}(\cdot)$ is the marginal distribution of the potential outcome $Y(z)\in\mathcal{Y}$ for $z=0,1$. Let $\{Y_{1}(1),Y_{1}(0)\}$ and $\{Y_{2}(1),Y_{2}(0)\}$ denote two independent replicates of $\{Y(1),Y(0)\}$, and $\tau_{h}$ can be re-expressed in a familiar form, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}]$. Two notable examples are the probability index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al., 2015), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al., 2011). These estimable quantities are alternative measurements for the population causal effect and can be straightforwardly interpreted. For example, the probability index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than $F_{0}$ and is robust against extreme values in the tails of the potential outcome distributions.
When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al., 1996), mortality event (Zhang et al., 2009) or dropout (Frangakis and Rubin, 1999), often have important implications even on the definition of a causal effect and require careful adjustment. The intermediate variable is not guaranteed to be balanced via randomization, and direct conditioning on the intermediate variable leads to sample selection bias and non-causal comparisons, even in randomized experiments. As a solution, the principal stratification framework (Frangakis and Rubin, 2002) segments the population into principal strata (subpopulations) according to joint potential values of the intermediate variable under alternative conditions. The stratum-specific causal effect—the principal causal effect—can thus be defined and is causally interpretable. However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation. Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function. The PGCE, under a linear contrast function, reduces to the usual principal average causal effect. As we demonstrate in due course, the PGCE can represent the stratum-specific probability index and strata-specific win ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations of interest.
Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al., 1996) and principal ignorability (Jo and Stuart, 2009; Ding and Lu, 2016), which are typically invoked for principal stratification analyses. The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, effectively restricting the number of potential strata. This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival, or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake. The principal ignorability assumption, on the other hand, requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the final outcome. For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al. (2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and further developed multiply robust and locally efficient estimators under the additional assumption of treatment unconfoundedness. In contrast, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
The contributions of this article are several-fold. First, we propose a class of PGCE estimands in the presence of an intermediate variable and develop complementary nonparametric identification formulas under treatment unconfoundedness, monotonicity, and principal ignorability. Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al., 1993, EIFs). Specifically, we show that when the contrast function $h(u,v)=u-v$, our EIFs reduce to those derived by Jiang et al. (2022) as a special case but are otherwise more general to accommodate arbitrary contrast functions. A major complexity in the context of a nonlinear contrast function, however, is that the estimating equations based on the EIFs require nuisance functions defined over pairs of observations. Thus, to obtain a feasible estimator, we define ratio $U$-statistics whose influence functions are the same as the EIFs. We show that the $U$-estimators are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions (including the treatment propensity score, principal score, and the potential outcome models) are correctly specified. When all nuisance functions are correctly specified, the $U$-estimators are locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds. The asymptotic properties of the $U$-estimators are developed upon the theory of empirical process (van der Vaart, 1998, §19), $U$-process (Arcones and Gine, 1993), as well as the theory of $U$-statistics (van der Vaart, 1998, §11-12). Finally, the construction of the multiply robust $U$-estimators facilitates the integration of machine learning estimators for the nuisance functions. To achieve full asymptotic efficiency, we learn each nuisance function using machine learning methods and operationalize the debiased machine learning $U$-estimators via cross-fitting (Chernozhukov et al., 2018). We discuss the modification needed for cross-fitting if a nonlinear contrast function is considered when defining the PGCE and prove the multiple robustness and efficiency property of the debiased machine learning estimators under mild regulation conditions (Chernozhukov et al., 2022; Escanciano and Terschuur, 2023).
The rest of this article is organized as follows. Section 2 introduces our causal estimands and identification assumptions for principal stratification. Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators. Section 4 derives the EIF for each PGCE estimand and develops the EIF-motivated feasible $U$-estimators and the debiased machine learning estimators. Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations. Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al., 2001), and Section 7 concludes. Detailed proofs of all theoretical results are available in the Online Supplement.
2 Notation, estimands and assumptions
2.1 Setting and notation
We consider a study (experimental or observational) with $n$ participants and two treatment conditions and let $Z_{i}$ denote the treatment assignment of participant $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 00 the control. We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively. The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$. Under the stable unit treatment value assumption (Rubin, 1980, SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$. Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d).
2.2 Principal generalized causal effects
We operate under the principal stratification framework (Frangakis and Rubin, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions. Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{11,10,01,00\}$. A key insight under the principal stratification framework is that although $D$ is a post-randomization variable, the stratum variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable. Therefore, based on any contrast function $h$, we define the principal generalized causal effect (PGCE) estimand as
$\tau_{h}^{s}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=S_{2}=s],~{}\text{for}~{}s\in\{11,10,01,00\},$ (2)
where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$ ($S_{1}=s$ and $S_{2}=s$ denote the strata membership for these two selected units). Of note, to ensure a proper causal interpretation of $\tau_{h}^{s}$, it is required that both elements of the pair for comparison are from the same principal stratum such that the comparison is made within a common subpopulation. If the two units are selected from different principal strata, the resulting quantity, $\tau_{h}^{s_{1},s_{2}}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$, remains estimable under this framework but generally lacks a meaningful causal interpretation; we thus do not further pursue this direction. Evidently, when a linear contrast function is specified as $h(u,v)=u-v$, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature; see, for example, Frangakis and Rubin (2002), Jiang et al. (2022) and Lipkovich et al. (2022). When a nonlinear contrast function is considered in defining the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum. The following two examples illustrate the PGCE with a specific contrast function.
Example 1 (Probability index under truncation by death)
When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status. That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment. Because the final outcome is oftentimes only well-defined without ambiguity when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al., 2009; Chen et al., 2024). In clinical research, Acion et al. (2006) argued that a more intuitive effect size can be measured by the probability index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$. Based on this contrast function, we can define the survivor probability index (SPI) as
$\mathrm{SPI}=\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=11\}.$ (3)
Compared to the SACE, the SPI is less sensitive to the skewness of the underlying distributions of the potential outcomes and heteroskedasticity between conditions and outliers. In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg, 1971).
Example 2 (Win estimands under noncompliance)
In experimental studies with ordinal outcomes, Wang and Pocock (2016) introduced the win ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$. This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{11,10,01,00\}$ refers to always-takers, compliers, defiers, and never-takers, respectively (Angrist et al., 1996). Therefore, the win ratio estimands can be extended to define the compliers win ratio (CWR) as
$\mathrm{CWR}=\frac{\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}}{\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}}.$ (4)
The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation. Alternatively, the complier win difference (CWD) or complier net benefit estimand on the difference scale is given as
$\mathrm{CWD}=\mathbb{P}\{Y_{1}(1)>Y_{2}(0)|S_{1}=S_{2}=10\}-\mathbb{P}\{Y_{1}(1)<Y_{2}(0)|S_{1}=S_{2}=10\}.$ (5)
The win estimands within other principal strata can be defined similarly. In the context of noncompliance, Mao (2024) defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE. Different from our approach, Mao (2024) used an instrumental variable approach and discussed point and interval identification of this estimand based on monotonicity and exclusion restriction.
2.3 Identification assumptions
Even in experimental studies, the PGCE estimands cannot be identified without additional assumptions because the pair of potential intermediate variables and that of potential final outcomes are never observed simultaneously. To identify the PGCE from the observed data, we require the following assumptions.
Assumption 1 (Treatment ignorability)
[…]
Assumption 1 rules out confounding between the treatment assignment and intermediate as well as final outcomes conditional on observed pre-treatment covariates. This assumption holds by design in randomized experiments. For observational studies, Assumption 1 can be satisfied if all confounders affecting the treatment assignment, the intermediate variable, and the final outcome are measured and included in the measured set of pre-treatment covariates, $X$.
Assumption 2 (Monotonicity)
$D_{i}(1)\geq D_{i}(0)$ for all $i$.
Assumption 2 states that the treatment assignment only affects the intermediate outcome in a non-negative (or non-positive if assume $D_{i}(1)\leq D_{i}(0)$) direction and restricts the number of possible strata by excluding the harmed or defiers stratum $S=01$ (or $S=10$). For example, when $D$ represents the survival status, monotonicity holds when the treatment is expected not to harm the survival of each unit. When $D$ represents actual treatment received in noncompliance problems, monotonicity holds if units in the control condition have no access to active treatment.
Assumption 3 (Mean principal ignorability)
Given the same values of pre-treatment covariates for a pair of units ($X_{1}=x_{1}$ and $X_{2}=x_{2}$), the expected contrast function is exchangeable across certain pairs of principal strata. That is,
$\begin{split}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2},X_{1}=x_{1},X_{2}=x_{2}]\\ =~{}&\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1}^{\prime},S_{2}=s_{2}^{\prime},X_{1}=x_{1},X_{2}=x_{2}]\end{split}$ (6)
for (a) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{10}=\{(10,10),(11,10),(10,00),(11,00)\}$; (b) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{11}=\{(11,11),(10,11)\}$; (c) $(s_{1},s_{2})\neq(s_{1}^{\prime},s_{2}^{\prime})\in\mathcal{S}_{00}=\{(00,00),(00,10)\}$.
Assumption 3 rules out unmeasured confounding between the principal strata membership and the final outcome, ensuring that the PGCE components can be directly estimated from the observed data. For example, under Assumptions 1-3, the conditional expectations in group $\mathcal{S}_{10}$ are all equal to $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$. In other words, within the observed data stratum defined by $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, conditional on pre-treatment covariates, the combinations of possible latent principal strata, $(S_{1}=10,S_{2}=10)$, $(S_{1}=11,S_{2}=10)$, $(S_{1}=10,S_{2}=00)$, and $(S_{1}=11,S_{2}=00)$, are exchangeable with respect to the expected contrast. Similarly, the conditional expected contrasts within group $\mathcal{S}_{11}$ and $\mathcal{S}_{00}$ are all equal to estimable conditional expectations, $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1,X_{1},X_{2}]$ and $\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0,X_{1},X_{2}]$, respectively. That is, with Assumption 3, each one of the subpopulations, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$, $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=1)$, and $(Z_{1}=1,D_{1}=0,Z_{2}=0,D_{2}=0)$, can be viewed as mixtures of its corresponding subgroups defined by observed covariates only.
Remark 1
[…]
Assumptions 2 and 3 are instrumental in achieving nonparametric identification of the PGCE; however, their applicability, in general, cannot be directly tested using observed data alone without prior knowledge. Although our primary focus is identifying and efficiently estimating the PGCE under these two structural assumptions, we provide a confounding function approach for sensitivity analyses under assumed violations of these two structural assumptions in Section S7 of the Online Supplement.
3 Nonparametric identification
3.1 Identification of the PGCE estimands
To obtain nonparametric identification formulas for the PGCE estimands, we start by defining components of the observed data distribution, i.e., the treatment probability, the principal score, and the mean of the outcome contrast conditional on covariates (referred to as pairwise outcome mean hereafter for brevity). We consider both the observational and experimental studies and define the treatment probability or the propensity score, by $\pi(X)=\mathbb{P}(Z=1|X)$, which is the probability of a participant receiving the treatment conditional on pre-treatment covariates (Rosenbaum and Rubin, 1983). In addition, the principal score, $e_{s}(X)=\mathbb{P}(S=s|X)$, is defined as the conditional probability that a participant belongs to the principal stratum $s$ given the covariates (Ding and Lu, 2016), for $s\in\{10,00,11\}$ under Assumption 2. Furthermore, we use averages over the distribution of covariates, $\pi=\mathbb{E}\{\pi(X)\}=\mathbb{P}(Z=1)$ and $e_{s}=\mathbb{E}\{e_{s}(X)\}=\mathbb{P}(S=s)$, to denote population fractions of receiving treatment $z$ or belonging to stratum $s$, respectively. Lastly, we define the pairwise outcome mean within the observed stratum $(Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2})$, as
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=z_{1},D_{1}=d_{1},Z_{2}=z_{2},D_{2}=d_{2},X_{1},X_{2}\},$
for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Below, we provide some examples of the pairwise outcome mean models for specific contrast functions.
Example 3 (Pairwise mean function for average treatment effect)
For studying the average treatment effect, the contrast function is specified as $h(u,v)=u-v$. If the potential outcomes are continuous, a Gaussian regression model is often assumed for the potential outcomes, i.e., $Y(z)|D,X\sim\mathcal{N}(f_{z}(D,X),\sigma^{2})$. Here, $f_{z}(D,X)$ is the mean function of $D$ and $X$ given treatment status $Z=z$. We then have the pairwise outcome mean function given by
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2}).$ (7)
Example 4 (Pairwise mean function for probability index)
For studying the probability index estimand, the contrast function is specified as $h(u,v)=\mathbbm{1}(u\geq v)$. If the potential outcomes are continuous, a Gaussian regression model can be considered to represent the mean potential outcomes as in Example 3. In this case, following Wolfe and Hogg (1971), one can derive the pairwise outcome mean as
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\Phi\left\{\frac{f_{z_{1}}(d_{1},X_{1})-f_{z_{2}}(d_{2},X_{2})}{\sqrt{2}\sigma}\right\}.$ (8)
Example 5 (Pairwise outcome mean for win estimands)
Consider the win estimands developed in Example 2, where the contrast function is given by the vector $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (the win estimands are then up to transformations of the two-dimensional estimand). For ordinal outcomes, if we specify the proportional odds or ordinal logistic regression model for $Y(z)$ conditional on $D$ and $X$, Mao (2017) has shown that the resulting pairwise outcome mean functions are given by the two-dimensional vector
$\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ (9)
=$\quad=\left\{\sum_{q=1}^{Q}\sum_{q^{\prime}=1}^{q-1}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2}),\sum_{q=1}^{Q}\sum_{q^{\prime}=q+1}^{Q}p_{z_{1}q}(d_{1},X_{1})p_{z_{2}q^{\prime}}(d_{2},X_{2})\right\},$
where $p_{zq}(D,X)$ is the conditional probability of $Y(z)=q\in\{1,\ldots,Q\}$ given $D$ and $X$, $Q$ is the largest category of the ordinal outcome.
For estimating the principal scores, we first define the conditional probability, $p_{z}(X)=\mathbb{P}(D=1|Z=z,X)$, with population fraction $p_{z}=\mathbb{E}\{p_{z}(X)\}=\mathbb{P}(D=1|Z=z)$. Then under Assumption 2, the principal scores can be expressed as (Jiang et al., 2022)
$e_{10}(X)=p_{1}(X)-p_{0}(X),~{}~{}~{}e_{00}(X)=1-p_{1}(X),~{}~{}~{}e_{11}(X)=p_{0}(X).$
This is because, with monotonicity, the observed stratum $(Z=0,D=1)$ only contains stratum $S=11$, and observed stratum $(Z=1,D=0)$ only contains stratum $S=00$. The remaining principal score for stratum $S=10$ can thus be obtained by subtracting $e_{11}(X)$ and $e_{00}(X)$ from one. The population fractions of the three principal strata can be similarly connected as $e_{10}=p_{1}-p_{0}$, $e_{00}=1-p_{1}$, and $e_{11}=p_{0}$. We will refer to $\{p_{0}(X),p_{1}(X)\}$ as principal scores due to the one-to-one mapping to $\{e_{10}(X),e_{00}(X),e_{11}(X)\}$. These elements then prepare us to state the set of nonparametric identification results in the following Theorem 1.
Theorem 1 (Nonparametric identification)
Suppose Assumptions 1-3 hold, $e_{s}>0$ for $s\in\{10,00,11\}$, and $10<\pi(X)<$ 1 for all $X\in\mathcal{X}$, the support of covariates. Then, based on different components of the observed data distribution, the following identification formulas for each PGCE can be obtained:
  • (a)
    Based on the treatment probability and the principal score,
    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}}{1-p_{1}}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{00}(X_{2})}{1-p_{1}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{e_{11}(X_{1})}{p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{D_{2}}{p_{0}}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}h(Y_{1},Y_{2})\right];$
  • (b)
    Based on the treatment probability and the pairwise outcome mean,
    $\tau_{h}^{10}$ $=\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
    ×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\mu_{1100}(X_{1},X_{2})\right),$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-D_{1}Z_{1}/\pi(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-D_{2}Z_{2}/\pi(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left(\left[\frac{D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{0}}\right]\left[\frac{D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{0}}\right]\mu_{1101}(X_{1},X_{2})\right);$
  • (c)
    Based on the principal score and the pairwise outcome mean,
    $\tau_{h}^{10}$ $=\mathbb{E}\left[\left\{\frac{p_{1}(X_{1})-p_{0}(X_{1})}{p_{1}-p_{0}}\right\}\left\{\frac{p_{1}(X_{2})-p_{0}(X_{2})}{p_{1}-p_{0}}\right\}\mu_{1100}(X_{1},X_{2})\right],$
    $\tau_{h}^{00}$ $=\mathbb{E}\left[\left\{\frac{1-p_{1}(X_{1})}{1-p_{1}}\right\}\left\{\frac{1-p_{1}(X_{2})}{1-p_{1}}\right\}\mu_{1000}(X_{1},X_{2})\right],$
    $\tau_{h}^{11}$ $=\mathbb{E}\left[\left\{\frac{p_{0}(X_{1})}{p_{0}}\right\}\left\{\frac{p_{0}(X_{2})}{p_{0}}\right\}\mu_{1101}(X_{1},X_{2})\right].$
We use $\tau_{h}^{10}$ as an example to elaborate on the intuitions behind Theorem 1, and the identification of the other two PGCEs follows the same strategy. In Theorem 1(a), $\tau_{h}^{10}$ is expressed as a doubly inverse probability-weighted average of the observed contrast function; the essential part of the weighting scheme leverages the principal scores to tilt all pairs of $(\mathcal{O}_{i},\mathcal{O}_{j})$ ($i\neq j$) in the observed stratum $(Z_{1}=1,D_{1}=1,Z_{2}=0,D_{2}=0)$ to represent the covariate distribution in the principal stratum $S=10$. Specifically, under Assumptions 1 and 2, we have
$\mathbb{E}\left[\left\{\frac{e_{10}(X_{1})}{p_{1}-p_{0}}\frac{D_{1}}{p_{1}(X_{1})}\frac{Z_{1}}{\pi(X_{1})}\right\}\left\{\frac{e_{10}(X_{2})}{p_{1}-p_{0}}\frac{1-D_{2}}{1-p_{0}(X_{2})}\frac{1-Z_{2}}{1-\pi(X_{2})}\right\}\bigg{\rvert}X_{1},X_{2}\right]=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}},$
which is proportional to the conditional probability of each pair belonging to the principal stratum $S=10$, and by Bayes’ Theorem, is equivalent to the density ratio of $(X_{1},X_{2})$ conditional and unconditional on $S_{1}=S_{2}=10$. This ensures the identification of $\tau_{h}^{10}$. In Theorem 1(b), while $\tau_{h}^{10}$ is expressed using the treatment probability and the pairwise outcome mean, it can be shown that the conditional expectation of the weight attached to the pairwise outcome mean, under Assumptions 1 and 2, is,
$\mathbb{E}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1})-D_{1}(1-Z_{1})/\{1-\pi(X_{1})\}}{p_{1}-p_{0}}\right]\right.$
×$\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2})-D_{2}(1-Z_{2})/\{1-\pi(X_{2})\}}{p_{1}-p_{0}}\right]\bigg{\rvert}X_{1},X_{2}\right)=\frac{e_{10}(X_{1})e_{10}(X_{2})}{e_{10}^{2}}.$
Therefore, $\tau_{h}^{10}$ is identified as the pairwise outcome mean function averaged over the density $f(X_{1},X_{2}|S_{1}=S_{2}=10)/f(X_{1},X_{2})$. Finally, in Theorem 1(c), $\tau_{h}^{10}$ is expressed using the principal score and pairwise outcome mean, where, in comparison with Theorem 1(b), the treatment probability and intermediate outcome, $DZ/\pi(X)-D(1-Z)/\{1-\pi(X)\}$, is replaced by the principal score, $p_{1}(X)-p_{0}(X)$. The conditional expectation of the weight on covariates attached to the pairwise outcome mean, under Assumptions 1 and 2, remains $e_{10}(X_{1})e_{10}(X_{2})/e_{10}^{2}$, and hence $\tau_{h}^{10}$ is identified as an average of the pairwise outcome mean function over the target covariate density.
Remark 2
Theorem 1 is a natural extension of the results for nonparametric identification of the principal average causal effects (Jiang et al., 2022, Theorem 1) to the PCGEs based on an arbitrary contrast function $h$; and it reduces to the Theorem 1 in Jiang et al. (2022) when $h(u,v)=u-v$ with $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})=\mu_{z_{1}d_{1}}(X_{1})-\mu_{z_{2}d_{2}}(X_{2})$.
3.2 Estimators based on nonparametric identification formulas
For point estimation, a common practice for operationalizing the nonparametric identification formulas is to assume parametric models for nuisance functions, i.e., $\pi(X;\alpha)$, $p_{z}(X;\beta)$ or $e_{s}(X;\beta)$, and $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma)$, where $\alpha$, $\beta$, and $\gamma$ are finite-dimensional nuisance parameters, with their corresponding true values denoted by $\alpha^{*}$, $\beta^{*}$, and $\gamma^{*}$. We write the maximum likelihood estimators for these finite-dimensional parameters as $\widehat{\alpha}$, $\widehat{\beta}$, and $\widehat{\gamma}$. To aid further discussions of statistical properties of estimators for the PGCEs, we introduce notations $\mathcal{M}_{\pi}$, $\mathcal{M}_{p}$, and $\mathcal{M}_{\mu}$ to denote conditions where the treatment probability, the principal score, and the outcome mean are correctly specified, respectively. Specifically, under $\mathcal{M}_{\pi}$, we have $\pi(X;\alpha^{*})=\pi(X)$; under $\mathcal{M}_{p}$, we have $p_{z}(X;\beta^{*})=p_{z}(X)$ and $e_{s}(X;\beta^{*})=e_{s}(X)$ for $z\in\{0,1\}$ and $s\in\{10,00,11\}$; and under $\mathcal{M}_{\mu}$, we have $\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2};\gamma^{*})=\mu_{z_{1}d_{1}z_{2}d_{2}}(X_{1},X_{2})$ for $z_{1},z_{2},d_{1},d_{2}\in\{0,1\}$. Additionally, we use ‘+’ in the subscript to denote conditions where more than one nuisance function is correctly specified, and $\cup$’ to denote the condition where at least one nuisance function is correctly specified.
We can obtain plug-in estimators by replacing components in the identification formulas with their empirical counterparts and expectations with empirical averages. For convenience, we first introduce the following notations: $\mathbb{E}_{n}f(\mathcal{O})=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ for any $f(\mathcal{O})$, and $\mathbb{U}_{n}f(\mathcal{O}_{1},\mathcal{O}_{2})=\binom{n}{2}^{-1}\sum_{i<j}f(\mathcal{O}_{i},\mathcal{O}_{j})$ for any $f(\mathcal{O}_{1},\mathcal{O}_{2})$. Along the lines in Jiang et al. (2022), we define the following quantity for any function $f(Y,D,X)$:
[…]
One example of (10) is having $f(Y_{z},D_{z},X)=D_{z}$, in which case we obtain the following function:
$\psi_{D_{z}}(\mathcal{O})=\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X)\}}{\mathbb{P}(Z=z|X)}+p_{z}(X).$ (11)
Equation (11) is useful in the estimation of $p_{z}$, as it leads to a doubly robust estimator once we treat the intermediate variable as the outcome (Tsiatis, 2006):
$\widehat{p}_{z}\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})=\mathbb{E}_{n}\left[\frac{\mathbbm{1}(Z=z)\{D-p_{z}(X;\widehat{\beta})\}}{\pi(X;\widehat{\alpha})}+p_{z}(X;\widehat{\beta})\right],$ (12)
which improves upon the plug-in estimator $\mathbb{E}_{n}p_{z}(X;\widehat{\beta})$ in terms of estimation efficiency and robustness, and is consistent for $p_{z}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$.
Plug-in estimators for the PGCEs take the form of $U$-statistics (van der Vaart, 1998). For example, Theorem 1(a) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated treatment probability and the principal score:
$\widehat{\tau}_{h,\pi+p}^{10}=\mathbb{U}_{n}\left[\left\{\frac{e_{10}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{D_{1}}{p_{1}(X_{1};\widehat{\beta})}\frac{Z_{1}}{\pi(X_{1};\widehat{\alpha})}\right\}\left\{\frac{e_{10}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\frac{1-D_{2}}{1-p_{1}(X_{2};\widehat{\beta})}\frac{1-Z_{2}}{1-\pi(X_{2};\widehat{\alpha})}\right\}h(Y_{1},Y_{2})\right].$
Theorem 1(b) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated propensity score and the pairwise outcome mean:
$\widehat{\tau}_{h,\pi+\mu}^{10}$ $=\mathbb{U}_{n}\left(\left[\frac{D_{1}Z_{1}/\pi(X_{1};\widehat{\alpha})-D_{1}(1-Z_{1})/\{1-\pi(X_{1};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\right.$
×$\qquad\qquad\times\left.\left[\frac{D_{2}Z_{2}/\pi(X_{2};\widehat{\alpha})-D_{2}(1-Z_{2})/\{1-\pi(X_{2};\widehat{\alpha})\}}{\widehat{p}_{1}-\widehat{p}_{0}}\right]\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right),$
and Theorem 1(c) motivates the following estimator for $\tau_{h}^{10}$ based on the estimated principal score and the pairwise outcome mean:
$\widehat{\tau}_{h,p+\mu}^{10}$ $=\mathbb{U}_{n}\left[\left\{\frac{p_{1}(X_{1};\widehat{\beta})-p_{0}(X_{1};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\left\{\frac{p_{1}(X_{2};\widehat{\beta})-p_{0}(X_{2};\widehat{\beta})}{\widehat{p}_{1}-\widehat{p}_{0}}\right\}\mu_{1100}(X_{1},X_{2};\widehat{\gamma})\right].$
The constructions of plug-in estimators for the remaining PGCEs are similar and, therefore, omitted. These estimators have stronger requirements on model specification to estimate the PGCEs consistently. Specifically, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment probability and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
Remark 3
The weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score, which is an extension of the inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable. The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately after applying the Taylor’s expansion on the ratio estimators (Jiang et al., 2022); see the Online Supplement for additional details. The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be straightforwardly obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2). The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).
4 Multiply robust and efficient $U$-estimators
Although implementing the plug-in estimators motivated by the nonparametric identification formulas is straightforward, their various forms and potential inefficiency call for alternative estimators with improved statistical properties. Here, we develop alternative $U$-estimators that exploit information across all nuisance functions, including the treatment assignment, the intermediate variable, and the final outcome of interest.
4.1 The efficient influence function for the PGCE
We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis, 2006; Mao, 2017). Second, the asymptotic variance lower bound is represented by the variance of the EIF, providing a useful benchmark to define efficient estimators. Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al., 2018; Rotnitzky et al., 2021). We will explore these properties by first formalizing the EIF for the PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
$\varphi_{h}^{s}(\mathcal{O})\equiv\frac{1}{\tau_{D}^{s}}\varphi_{h,N}^{s}(\mathcal{O})-\frac{\tau_{h,N}^{s}}{\tau_{D}^{s}{}^{2}}\varphi_{D}^{s}(\mathcal{O})=\frac{\varphi_{h,N}^{s}(\mathcal{O})-\tau_{h}^{s}\varphi_{D}^{s}(\mathcal{O})}{\tau_{D}^{s}},$
where $\varphi_{h,N}^{s}(\mathcal{O})$ and $\varphi_{D}^{s}(\mathcal{O})$ are EIFs for $\tau_{h,N}^{s}$ and $\tau_{D}^{s}$, respectively. Let $F_{zd}(\cdot|x)$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},X_{i})=(z,d,x)$, and $F_{X}(\cdot)$ denote the marginal distribution of $X$. We can further define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(x)h(Y,y)\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(x)h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$. Then, the following results are obtained.
Theorem 2 (EIF)
Under Assumptions 1-3, (a) the EIF for $\tau_{h}^{10}$ is $\varphi_{h}^{10}(\mathcal{O})=\{\varphi_{h,N}^{10}(\mathcal{O})-\tau_{h}^{10}\varphi_{D}^{10}(\mathcal{O})\}/(p_{1}-p_{0})^{2}$, where $\varphi_{h,N}^{10}(\mathcal{O})=\phi_{h}^{10}(\mathcal{O})-2\tau_{h,N}^{10}$ and
$\phi_{h}^{10}(\mathcal{O})=$ $~{}\frac{ZDe_{10}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,00}^{10}(Y)-m_{1}^{10}(X)\right\}+\frac{(1-Z)(1-D)e_{10}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,11}^{10}(Y)-m_{0}^{10}(X)\right\}$
+$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
with $m_{1}^{10}(X)=\int e_{10}(x)\mu_{1100}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{10}(X)=\int e_{10}(x)\mu_{1100}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{10}(\mathcal{O})=2\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}(p_{1}-p_{0})-2\tau_{D}^{10}.$
(b) The EIF for $\tau_{h}^{00}$ is $\varphi_{h}^{00}(\mathcal{O})=\{\varphi_{h,N}^{00}(\mathcal{O})-\tau_{h}^{00}\varphi_{D}^{00}(\mathcal{O})\}/(1-p_{1})^{2}$, where $\varphi_{h,N}^{00}(\mathcal{O})=\phi_{h}^{00}(\mathcal{O})-2\tau_{h,N}^{00}$ and
$\phi_{h}^{00}(\mathcal{O})=$ $~{}\frac{Z(1-D)}{\pi(X)}\left\{\nu_{1,00}^{00}(Y)-m_{1}^{00}(X)\right\}+\frac{(1-Z)(1-D)e_{00}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,10}^{00}(Y)-m_{0}^{00}(X)\right\}$
$+\psi_{1-D_{1}}(\mathcal{O})\left\{m_{1}^{00}(X)+m_{0}^{00}(X)\right\},$
with $m_{1}^{00}(X)=\int e_{00}(x)\mu_{1000}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{00}(X)=\int e_{00}(x)\mu_{1000}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{00}(\mathcal{O})=2\psi_{1-D_{1}}(\mathcal{O})(1-p_{1})-2\tau_{D}^{00}.$
(c) The EIF for $\tau_{h}^{11}$ is $\varphi_{h}^{11}(\mathcal{O})=\{\varphi_{h,N}^{11}(\mathcal{O})-\tau_{h}^{11}\varphi_{D}^{11}(\mathcal{O})\}/p_{0}^{2}$, where $\varphi_{h,N}^{11}(\mathcal{O})=\phi_{h}^{11}(\mathcal{O})-2\tau_{h,N}^{11}$ and
$\phi_{h}^{11}(\mathcal{O})=$ $~{}\frac{ZDe_{11}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,01}^{11}(Y)-m_{1}^{11}(X)\right\}+\frac{(1-Z)D}{1-\pi(X)}\left\{\nu_{0,11}^{11}(Y)-m_{0}^{11}(X)\right\}$
$+\psi_{D_{0}}(\mathcal{O})\left\{m_{1}^{11}(X)+m_{0}^{11}(X)\right\},$
with $m_{1}^{11}(X)=\int e_{11}(x)\mu_{1101}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{11}(X)=\int e_{11}(x)\mu_{1101}(x,X)\mathrm{d}F_{X}(x)$, and
$\varphi_{D}^{11}(\mathcal{O})=2\psi_{D_{0}}(\mathcal{O})p_{0}-2\tau_{D}^{11}.$
The EIFs in Theorem 2 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al. (2022) with respect to the distribution of observed data, $\mathcal{O}$. The derivation for $\varphi_{D}^{s}(\mathcal{O})$ then follows the chain rule. For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
$\tau_{h}^{s}=\mathbb{E}\{\mathfrak{h}_{1}^{s}(Y)|Z=1\}+\mathbb{E}\{\mathfrak{h}_{0}^{s}(Y)|Z=0\},$
where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$. Importantly, the EIF for the PGCE remains unchanged after this re-expression since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged. Based on Theorem 2, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
Remark 4
Without considering the intermediate variable, the EIF for the generalized causal effect estimand, $\tau_{h}$ in (1), $\varphi_{h}(\mathcal{O})$, can be derived using the same approach as $\varphi_{h}(\mathcal{O})=\phi_{h}(\mathcal{O})-2\tau_{h}$, where
$\phi_{h}(\mathcal{O})=\frac{Z}{\pi(X)}\left\{\nu_{1}(Y)-m_{1}(X)\right\}+\frac{1-Z}{1-\pi(X)}\left\{\nu_{0}(Y)-m_{0}(X)\right\}+m_{1}(X)+m_{0}(X),$
with $\nu_{z}(Y)=\int h(Y,y)\mathrm{d}F_{z}(y|x)\mathrm{d}F_{X}(x)$, $m_{1}(X)=\int\mu_{10}(X,x)\mathrm{d}F_{X}(x)$ and $m_{0}(X)=\int\mu_{10}(x,X)\allowbreak\mathrm{d}F_{X}(x)$. The function $\mu_{10}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=1,Z_{2}=0,X_{1},X_{2}\}$ is the pairwise outcome mean within the observed stratum $(Z_{1}=1,Z_{2}=0)$. This coincides with the EIF developed in Mao (2017) using the projection approach (Tsiatis, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
4.2 Multiply robust estimators
The EIFs given in Theorem 2 imply the semiparametric efficiency lower bound and pave the way for formulating estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator. We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 2, thereby achieving semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that each PGCE can be rewritten as the following:
$\tau_{h}^{10}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{10})}{\{\mathbb{E}(\psi_{D_{1}}-\psi_{D_{0}})\}^{2}},~{}~{}~{}\tau_{h}^{00}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{00})}{\{\mathbb{E}(\psi_{1-D_{1}})\}^{2}},~{}~{}~{}\tau_{h}^{11}=\frac{2^{-1}\mathbb{E}(\phi_{h}^{11})}{\{\mathbb{E}(\psi_{D_{0}})\}^{2}}.$ (13)
Continuing with parametric models for nuisance functions, the denominators in (13) can be estimated using the doubly robust estimator given in (12). Specifically, let $\theta$ denote the collection of parametric nuisance function parameters and $\vartheta$ denote a subset of $\theta$, i.e., $\theta=\{\alpha,\beta,\gamma\}$ and $\vartheta=\{\alpha,\beta\}$, with corresponding maximum likelihood estimators $\widehat{\theta}=\{\widehat{\alpha},\widehat{\beta},\widehat{\gamma}\}$ and $\widehat{\vartheta}=\{\widehat{\alpha},\widehat{\beta}\}$, as well as true values $\theta^{*}=\{\alpha^{*},\beta^{*},\gamma^{*}\}$ and $\vartheta^{*}=\{\alpha^{*},\beta^{*}\}$. Then, for example, the denominator of $\tau_{h}^{10}$ given in (13) can be estimated by $[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}$, where $\mathbb{E}_{n}\psi_{D_{z}}(\widehat{\vartheta})\equiv\mathbb{E}_{n}\psi_{D_{z}}(\mathcal{O};\widehat{\alpha},\widehat{\beta})$ defined in (12).
Estimating the numerators in (13) is more complicated. To target these terms, we construct the following functions of the observed pairs:
$g_{h}^{10}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
$=2^{-1}\bigg{[}\frac{Z_{i}D_{i}e_{10}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})(1-D_{j})e_{10}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1100}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}D_{j}e_{10}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})(1-D_{i})e_{10}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1100}(X_{j},X_{i};\gamma)\right\}$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
+$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg{]},$
and
$g_{h}^{00}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$
$=2^{-1}\bigg{[}\frac{Z_{i}(1-D_{i})}{\pi(X_{i};\alpha)}\times\frac{(1-Z_{j})(1-D_{j})e_{00}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1000}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}(1-D_{j})}{\pi(X_{j};\alpha)}\times\frac{(1-Z_{i})(1-D_{i})e_{00}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1000}(X_{j},X_{i};\gamma)\right\}$
$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1000}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\psi_{1-D_{1}}(\mathcal{O}_{j};\alpha,\beta)\psi_{1-D_{1}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1000}(X_{j},X_{i};\gamma)\bigg{]},$
and
$g_{h}^{11}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ $=2^{-1}\bigg{[}\frac{Z_{i}D_{i}e_{11}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})D_{j}}{1-\pi(X_{j};\alpha)}\left\{h(Y_{i},Y_{j})-\mu_{1101}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}D_{j}e_{11}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})D_{i}}{1-\pi(X_{i};\alpha)}\left\{h(Y_{j},Y_{i})-\mu_{1101}(X_{j},X_{i};\gamma)\right\}$
$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\mu_{1101}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\mu_{1101}(X_{j},X_{i};\gamma)\bigg{]},$
for $1\leq i\neq j\leq n$. For $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$, it is straightforward to show that $2\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|\mathcal{O}_{i}\}=\phi_{h}^{s}(\mathcal{O}_{i})$ and $\mathbb{E}\{g_{h}^{s}(\mathcal{O}_{i},\allowbreak\mathcal{O}_{j};\theta^{*})\}=2^{-1}\mathbb{E}\{\phi_{h}^{s}(\mathcal{O}_{i})\}=\tau_{h,N}^{s}$. Thus, the numerators can be estimated using an empirical average of all pairs as $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})\equiv\binom{n}{2}^{-1}\sum_{i<j}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta})$ for $s\in\{10,00,11\}$. Therefore, we can obtain the following estimators for the PGCEs expressed in (13):
$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (14)
which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 3.
Theorem 3 (Multiply robustness and efficiency under parametric working models)
Under Assumptions 1-3, and for some $\varepsilon\in(0,1)$, $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, the estimator $\widehat{\tau}_{h}^{s}$ in (14) is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
Under Theorem 3, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified. The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator. The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$. The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions. The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao (2017) and Theorem 3 of Jiang et al. (2022). As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator. Details of the proof are provided in Section S4.2 of the Online Supplement. Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
4.3 Debiased machine learning estimators
Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies and is often built upon the orthogonal moment functions (Chernozhukov et al., 2018, 2022) or the EIFs. Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al., 2021). Compared to parametric working models, leveraging machine learning working models for the nuisance functions is more advantageous for its robustness against potential misspecification. To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al., 2018). Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2023). Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition =$\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq N\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov et al. (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$. The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
Example 6 (Sample splitting for $U$-statistics)
Consider a scenario where $n=10$ and $K=3$. We can first form the partition =$\mathcal{I}=\{I_{1},I_{2},I_{3}\}$ with $I_{1}=\{1,2,3\}$, $I_{2}=\{4,5,6,7\}$, and $I_{3}=\{8,9,10\}$. Then, the partition $\mathcal{P}$ is formed as
$\{(I_{1},I_{1}),(I_{1},I_{2}),(I_{1},I_{3}),(I_{2},I_{2}),(I_{2},I_{3}),(I_{3},I_{3})\}.$
For $\mathcal{P}$, elements formed by $I_{k_{1}}$ and $I_{k_{2}}$ with $k_{1}<k_{2}$ are more straightforward as the Cartesian product of $I_{k_{1}}$ and $I_{k_{2}}$. For instance,
[…]
Elements formed by $I_{k}$ itself, on the other hand, need to exclude pairs consisting of the same indices, e.g., $(I_{1},I_{1})=\{(1,2),(1,3),(2,3)\}$. According to this sampling-splitting scheme, all $\binom{10}{2}=45$ possible pairs are divided into $L=6$ subsets.
With a slight abuse of notation, let $\widehat{\theta}=\{\widehat{\pi},\widehat{p}_{z},\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\widehat{\vartheta}=\{\widehat{\pi},\widehat{p}_{z}\}$ denote machine learning estimators of nuisance functions $\theta=\{\pi,p_{z},\mu_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\vartheta=\{\pi,p_{z}\}$, respectively, as well as true functions $\theta^{*}=\{\pi^{*},p_{z}^{*},\mu_{z_{1}d_{1}z_{2}d_{2}}^{*}\}$ and $\vartheta^{*}=\{\pi^{*},p_{z}^{*}\}$. For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$). By pooling these intermediate estimates together, we can obtain the following cross-fitting estimators for the PGCEs:
$\widehat{\tau}_{h}^{10}=\frac{\mathbb{U}_{n,L}g_{h}^{10}(\widehat{\theta})}{[\mathbb{E}_{n,K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\mathbb{U}_{n,L}g_{h}^{00}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\mathbb{U}_{n,L}g_{h}^{11}(\widehat{\theta})}{\{\mathbb{E}_{n,K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$ (15)
where
$\mathbb{U}_{n,L}g_{h}^{s}(\widehat{\theta})=\binom{n}{2}^{-1}\sum_{l=1}^{L}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})$
and
$\mathbb{E}_{n,K}\psi_{D_{z}}(\widehat{\vartheta})=n^{-1}\sum_{k=1}^{K}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}).$
Remark 5
Estimators given in (15) are the type II estimator proposed in Chernozhukov et al. (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
$\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})=|P_{l}|^{-1}\sum_{(i,j)\in P_{l}}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\hat{\vartheta}_{k})=|I_{k}|^{-1}\sum_{i\in I_{k}}\psi_{D_{z}}(\mathcal{O}_{i};\widehat{\vartheta}_{k}),$
on $P_{l}$ and $I_{k}$, respectively, where $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ are correspondingly estimated on $P_{l}^{c}$ and $I_{k}^{c}$. Then, define
$\overline{\mathbb{U}}_{L}g_{h}^{s}(\widehat{\theta})=L^{-1}\sum_{l=1}^{L}\mathbb{U}_{|P_{l}|}g_{h}^{s}(\widehat{\theta}_{l})~{}~{}~{}\text{and}~{}~{}~{}\overline{\mathbb{E}}_{K}\psi_{D_{z}}(\widehat{\vartheta})=K^{-1}\sum_{k=1}^{K}\mathbb{E}_{|I_{k}|}\psi_{D_{z}}(\widehat{\vartheta}_{k}),$
and the estimator for the PCEs can be formed as
$\widehat{\tau}_{h}^{10}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{10}(\widehat{\theta})}{[\overline{\mathbb{E}}_{K}\{\psi_{D_{1}}(\widehat{\vartheta})-\psi_{D_{0}}(\widehat{\vartheta})\}]^{2}},~{}~{}~{}\widehat{\tau}_{h}^{00}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{00}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{1-D_{1}}(\widehat{\vartheta})\}^{2}},~{}~{}~{}\widehat{\tau}_{h}^{11}=\frac{\overline{\mathbb{U}}_{L}g_{h}^{11}(\widehat{\theta})}{\{\overline{\mathbb{E}}_{K}\psi_{D_{0}}(\widehat{\vartheta})\}^{2}},$
which are the type I debiased machine learning estimators. We focus on the type II estimator because it is generally recommended for its numerical stability in pooling intermediate estimates based on split data, especially in smaller samples.
We now discuss the asymptotic properties of the debiased machine learning estimators in (15). Assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability. In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$. The following result summarizes the efficiency properties of (15).
Theorem 4 (Efficiency under nonparametric working models)
Under Assumptions 1-3, and
  • (a)
    $\mathbb{E}|g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})|^{2}<\infty$ and $\mathbb{E}|\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})|^{2}<\infty$;
  • (b)
[…]
  • (c)
[…]
  • (d)
    $g_{h,1}^{s}(\mathcal{O};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O},o;\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 2.
Theorem 4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al. (2022, Lemma 15) and Escanciano and Terschuur (2023, Lemma 4). To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$. Regulatory conditions in Theorem 4 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018). The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator, which has been commonly adopted in debiased machine learning literature (Vansteelandt and Dukes, 2022; Wang et al., 2024). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.
5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 1, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15). In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
5.1 The probability index estimands with an intermediate variable
[…]
$D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\},$
$Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1),$
for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$. In this simulation study, we consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance), and assume the interest lies in estimating all three principal probability index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$). In addition, we create transformed covariates, $\widetilde{X}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\prime}$, where
$\widetilde{X}_{i1}=\exp(0.5X_{i1}),~{}\widetilde{X}_{i2}=\frac{X_{i2}}{1+X_{i1}},~{}\widetilde{X}_{i3}=\left(\frac{X_{i2}X_{i3}}{25}+0.6\right)^{3},~{}\widetilde{X}_{i4}=\frac{X_{i3}X_{i4}}{\sqrt{2}},$
to simulate scenarios where the nuisance functions are misspecified. Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise. Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries. We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
Figure 1 presents the results of the five estimators under comparison in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions). Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure 1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In comparison to the plug-in moment estimators, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared to the triply robust estimator, the debiased machine learning estimator leads to a slight improvement and is generally less biased.
(a) $\tau_{h}^{10}$
(b) $\tau_{h}^{00}$
(c) $\tau_{h}^{11}$
Figure 1: Violin plots of estimators for the principal probability index estimand in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
5.2 The win ratio estimands with an intermediate variable
In the second simulation study, we consider the case of the win ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model:
$\log\frac{\mathbb{P}(Y_{i}(Z_{i})\leq q)}{\mathbb{P}(Y_{i}(Z_{i})>q)}=\eta_{q,0}+2Z_{i}-D_{i}+X_{i1}-X_{i2}+1.2X_{i3}-0.8X_{i4},$
for $q\in\{1,2\}$, where $\eta_{1,0}=-1$ and $\eta_{2,0}=1$. Transformed covariates are created following the procedure described in the previous simulation study. For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are similarly estimated via logistic regression. For the debiased machine learning estimator, we used the multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using the SuperLearner package. The debiased machine learning estimator adopts a five-fold cross-fitting ($K=5$ and $L=15$) procedure. We set the number of observations $n=2,000$ with $1,000$ replications. Violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure 2 in Section S6 of the Online Supplement.
From Figure 2, we observe that, similar to the estimation of the probability index estimand in Section 5.1, the plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified, while the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified. When at least two of the three nuisance functions are misspecified (shown in Web Figure 2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the rest in terms of estimation accuracy (bias).
(a) $\tau_{h}^{10}$
(b) $\tau_{h}^{00}$
(c) $\tau_{h}^{11}$
Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24. The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later. Follow-up surveys were administered to both groups 24, 36, and 48 months after the random assignment to collect data on the outcomes of interest. As in many open-label randomized controlled trials, substantial noncompliance to treatment assignment as an intermediate outcome occurred 1 year after randomization (15.4% [$857/5,577$] of participants assigned to treatment did not adhere to the assignment, and 50.6% [$1,854/3,663$] assigned to control switched to the treatment condition).
The proposed methods were implemented to examine the causal effect of Job Corps intervention, where the weekly earnings one year after the assignment were selected as the final outcome. As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates, including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
6.1 The SPI for continuous outcomes
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al. (2009). Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
The estimated results for SPI are given in Table 1, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. We observe that, for the always-employed (72.9% of all participants by estimation), the probability index is estimated as $0.539$ using the triply robust and the debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates under the plug-in moment methods are similar to the triply robust and debiased machine learning methods, which likely suggests that each parametric working model may be approximately correctly specified.
Table 1: Estimated SPI for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval (CI) is given in the following parenthesis.
Methods SPI estimates
M1 0.521 (0.508, 0.533)
M2 0.518 (0.505, 0.530)
M3 0.525 (0.513, 0.538)
TR 0.539 (0.524, 0.555)
DML 0.539 (0.527, 0.552)
6.2 Principal win ratio for ordinal outcomes
We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance. In this example, we use the actual treatment status as an intermediate variable, similar to Chen and Flores (2015). Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$). To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al. (2018). Under this setting, we estimate the principal win ratios within each principal strata similar to the definition in (4). For methods based on parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are estimated via logistic regression. For the debiased machine learning approach, we used the multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using Super Learning; the estimator is fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
The estimated results are given in Table 2, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. The estimated proportions of always-takers, compliers, and never-takers in all participants are 50.1%, 34.0%, and 15.9%, respectively. The Jobs Corps intervention assignment positively affects participants’ earnings one year after the assignment for compliers and always-takers, and the effect appears to be more pronounced for compliers. For example, using the debiased machine learning estimator, there is a 27% increase in the chances of being in a higher earning category for a complier under treatment than a complier under control. The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1. For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win ratio is estimated to be closer to 1, and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1. Interestingly, in this reanalysis using a win ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable. In particular, the statistically significant effect among the always-takers is indicative of a positive psychological effect. That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence having a large chance to move into a higher earning category) compared to those who switched from their initial control assignment.
Table 2: Estimated win ratios within different compliance strata for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval is given in the following parenthesis.
Strata
Methods $10$ $00$ $11$
M1 1.216 (1.128, 1.304) 1.096 (0.955, 1.238) 1.117 (1.016, 1.218)
M2 1.206 (1.122, 1.289) 1.087 (1.025, 1.149) 1.096 (1.025, 1.168)
M3 1.204 (1.121, 1.286) 1.089 (1.025, 1.152) 1.096 (1.025, 1.167)
TR 1.197 (1.112, 1.282) 1.026 (0.920, 1.133) 1.131 (1.040, 1.223)
DML 1.270 (1.146, 1.393) 1.041 (0.940, 1.142) 1.167 (1.082, 1.252)
7 Discussion
In this article, we developed principal stratification methods for addressing a general class of PGCE estimands defined with a binary intermediate variable. Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-estimators that combine several different nuisance working functions. We showed that the $U$-estimators are triply robust in that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified. The $U$-estimators become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or by machine learning models with cross-fitting. In light of the increasing interest in pairwise comparison estimands in studies with a composite endpoint (Buyse, 2010; Bebu and Lachin, 2015) in biomedical research, our methods can also be useful when there is a need to additionally address a binary intercurrent event, such as treatment noncompliance or death (Bornkamp et al., 2021; Lipkovich et al., 2022; Mao, 2024)
The estimands and efficient $U$-estimators were developed to address a binary treatment and a binary intermediate outcome specifically but without restrictions on the final outcome. Although this is a standard setup for principal stratification analysis, additional work is needed to expand the identification conditions and results to address extensions of the PGCE estimands with a continuous treatment and/or a continuous intermediate variable (Antonelli et al., 2023). With continuous treatments, the definition of pairwise comparison estimands becomes non-trivial as there are infinitely many contrast functions (based on an infinitesimal increment of the treatment value) that one can specify to define $\tau_{h}$ even in the absence of an intermediate variable. On the other hand, with a continuous intermediate variable, there are infinitely many principal strata that can lead to conceptual and implementation challenges (Schwartz et al., 2011). Focusing on a linear contrast function, Lu et al. (2023) have developed nonparametric identification and semiparametric efficient estimators for principal causal effects with a continuous intermediate variable, which may inspire future theoretical development that allows for a more general nonlinear contrast function in the estimands definition.
Finally, as in all principal stratification analyses that achieve point identification, our $U$-estimators for estimating PGCEs depend on the monotonicity and the mean principal ignorability assumptions that are unverifiable from the observed data alone. In Section S7 of the Online Supplement, we have described a sensitivity function approach to obtain the bias-corrected $U$-estimators for estimating the PGCEs when either the monotonicity assumption or the mean principal ignorability assumption is violated. The bias-corrected $U$-estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed departures from the key identification assumptions.
Acknowledgement
Research in this article was partially supported by the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (NHLBI, grant number R00-HL141678). All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH.